OpenStudy (anonymous):

In how many ways can you write 18 as the sum of three counting numbers?

OpenStudy (zzr0ck3r):

hmm what class is this for?

OpenStudy (anonymous):

algebra 2/Trig

OpenStudy (zzr0ck3r):

Is this a homework problem, or are you just curious? This is a question for combinatorics.

OpenStudy (anonymous):

this is review for the test tomorrow. I think my teacher is trying to start us on pre calc for next year so i think thats what this is.

OpenStudy (zzr0ck3r):

Do you have any idea how you are supposed to solve this?

OpenStudy (anonymous):

on these type of combination problems we've been using factorials, and sometimes exponential functions, but I don't know where to start on this one

OpenStudy (zzr0ck3r):

3! = 6 6 + 6 + 6 10+7+1 10+5+3 10+6+2 .......

OpenStudy (zzr0ck3r):

are counting numbers just natural?

OpenStudy (anonymous):

so its 3 ! ?

OpenStudy (zzr0ck3r):

I can list much more than three if I understand the question right. But I don't think I do

OpenStudy (zzr0ck3r):

err much more than 6

OpenStudy (anonymous):

we've also been using pascals triangle if that helps.

OpenStudy (zzr0ck3r):

@Luis_Rivera can you explain please

OpenStudy (anonymous):

is it possible that it has something to do with the 18th row? or 18!/(3!15!) or something?

OpenStudy (zzr0ck3r):

I have no idea, maybe luis will answer.

OpenStudy (anonymous):

haha I hope so

OpenStudy (zzr0ck3r):

but I can list way more than 6 ways of writing 18 with the sum of three natural numbers

OpenStudy (anonymous):

10 + 5 + 3= 5 *2 + 5 + 3 primez : 5, 3, and 2 6 + 4 + 8 = 3*2 + 2*2 + 2^3 And on and on, notice that even though you see a 5, the actual numbers used total 3. Anyway that's how I was taught in ancient India

OpenStudy (zzr0ck3r):

ahh I see.

OpenStudy (zzr0ck3r):

so its how many numbers are used, not 3 numbers added together.

OpenStudy (zzr0ck3r):

bah im still confused...lol gl @liv2dance

OpenStudy (tkhunny):

We seem to be ignoring repeats. it does NOT say 3 DISTINCT values. Try expanding this \(\left(\sum\limits_{n = 1}^{16}x^{n}\right)^{3}\), and just reading the coefficient off the resulting \(x^{18}\) term. I get 136.

OpenStudy (anonymous):

I know it does not say distinct values, but they must be assumed, like in permutations of a word. Even though they do not say distinct you will not use the same word twice

OpenStudy (anonymous):

i think @tkhunny is right though.

OpenStudy (anonymous):

I don't see what the prime numbers have to do with finding the amount of (__+__+__) that equals 18

OpenStudy (tkhunny):

Generating function can be tedious, to be sure, but the idea is not overwhelming.

OpenStudy (anonymous):

Haha sorry. I'm almost positive @tkhunny is right. Thanks for everyones help though